LHC 2010 - latest news

Week 26: Machine Coordinators: Mike Lamont, Ralph Assmann


Latest plan: link PDF link XLS (version of 4 July, 13:50)

List of End of Fill studies (in work)

List of MD requests (in work)

Number of accesses: 6
Number of ramps: 11
Number of fills declared for physics: 6

Fastest turnaround: 4h48 (end stable beams to start stable beams)

Beam dumps on 5 ramps not making it into physics: 1) Losses from WS above threshold; 2) LSA software problems; 3) QPS trip; 4) Single collimator not ramping; 5) No longitudinal blowup during ramp;

Reason for end of 6 physics fills: 1) Electrical glitch x 2; 2) Instability-induced losses; 3) End of fill study x 2; 4) Internal beam dump interlock;

Day Short summary/plan Link to reports from 8:30 and 17:00 meetings
Monday July 5th
  • Stable beams continued...
Report from 8h30 meeting
Sunday July 4th
  • 00h46: Stable beams.
  • 01h35: Beam dump due to trip of RD1.LR1 from a glitch on the electrical network. Ramp down and precycle.
  • 03h21: Injection.
  • 04h23: Ramp started.
  • 05h09: Flat top.
  • 06h23: Stable beams.
  • 15h21: Adjust mode for switch on of transverse damper.
  • 16h34: Stable beams. Transverse damper is on. Vernier scans.
  • Beam uptime (until 14h45): 13h30
  • Stable beams (until 14h45): 9h51
Report from 8h30 meeting
Saturday July 3rd
  • 1h23: Threshold increased. Recycling
    Could not switch ON-STANDBY RB.A81. Calling TE/EPC piquet.
  • 3h04: Access PM85.
  • 5h20: Problem solved. Coming from 18 kV.
  • 5h26: QPS problem 6-7.
  • 6h40: Precycling.
  • 8h22: Injection.
  • 9h20: Start of ramp. 7 x 0.9e11 p per beam.
  • 10h05: Beam dump due to emittance measurement (wire scan) at 3 TeV: BLM threshold exceeded. Ramp down and precycle. Selected BLM thresholds increased by using monitor factor (BLM team).
  • 11h52: Loss of QPS OK on RCO in S56 and QQT13.L6B2/B1.
  • 12h29: RQT13.L6 and R6 trip for both beams. RCD and RCO in S56 and 67 trip while trying to sort out the QPS issue.
  • 12h34: Recycling affected circuits.
  • 13h29: Injection.
  • 13h59: Trip of sector 8-1.
  • Beam uptime (until 14h30): 2h10
  • 14h05: Access. QPS board changed for RB11.R8.
  • 18h09: Injection. Work on transverse emittance blow-up.
  • 20h26: FMCM trip on the MSD due to glitch.
  • 20h51: Injection. Work on transverse emittance blow-up.
  • 22h27: Ramp started.
  • 23h16: Flat top.
  • Beam uptime: 7h55
  • Stable beam: 0h00
Report from 8h30 meeting
Friday July 2nd
  • 1h31: Injection.
  • 3h04: Ramp started.
  • 3h57: Flat top. Issues with closure of separation bumps.
  • 7h32: Beam dump.
  • 9h43: Loss response checks in IR3.BLM response for injected beam hitting TCP and TCLA. Lost QPS OK on RD2.R5. Calling MP3. Masked.
  • 10h34: Prepare for access in IR5: RD2.R5 QPS.
  • 12h18: Access finished. Precycling.
  • 14h01: Beams injected. 7 times 1e11 p per beam.
  • 15h17: Start of ramp.
  • 16h06: Flat top.
  • 17h30: Stable beams.
  • 18h13: Beam dump with beam-beam induced instabilities. Rampdown and precycle. BLM intervention to change a card in SR6 on surface.
    Problem with RSD2.A81B1, the QPS has lost communication. RCO.A56B1 tripped during the ramp down.
  • 20h30: Injection.
  • 22h12: Ramp started.
  • 22h52: Beams dumped in ramp, due to loss of sector 81. Ramp down and precycle.
  • Beam uptime: 12h30
  • Stable beam: 0h43
Report from 8h30 meeting
Thursday July 1st
  • 02h56: Stable beams.
  • 07h30: Adjust mode. Collimators moved.
  • 07h41: Stable beams.
  • 10h38: Adjust mode. ALICE spearation changed to adjust luminosity.
  • 11h51: Setup of TCT's in IR2 for new separation orbit.
  • 14h53: Beams dumped with loss map by crossing 1/3 resonance.
  • 15h33: ATLAS access. LHCb access.
  • 19h13: Injection.
  • 20h31: Ramp started. Beams dumped with collimator interlock. Recover. Recyle. PS problems.
  • Beam uptime: 17h00
  • Stable beam: 7h31
Report from 8h30 meeting
Wednesday June 30th
  • 00h23: Main Power Supply failure in SPS.
  • 04h46: SPS recovered. Beam injection into LHC.
  • 05h22: Start ramp with 3e11 p per beam. 15% of beam 2 lost during ramp (chromaticity). At 3.5 TeV losses for beam 1 when collapsing separation.
  • 08h15: Stable beams.
  • 10h37: Beam dump. Voltage divider problem for dilution kicker. Possible to continue after beam dump expert masked some redundant channels. However, a GTO is broken and will require 1/2 day access at some point. For the moment operation can continue.
  • 10h47: Rampdown and precycle.
  • 12h52: Injection setup.
  • 13h49: Transverse damper studies.
  • 15h08: BPMS studies.
  • 15h21: Thunderstorm. Trip of sectors 12, 23, 34, 45, 56 + LHCb and ALICE dipoles + collimators. Recovery and precycle.
  • 20h45: Beam back. No QPS-OK from RCBXV2.R2. Lost QPS-OK from RB.A67, RQD/RQF.A67. Kick response measurements.
  • 23h01: Ramp started.
  • Beam uptime: 10h30
  • Stable beams: 2h22

Report from 8h30 meeting

Report at the LMC

Tuesday June 29th
  • 00h04: Ramp started. Longitudinal blowup not active.
  • 01h00: Requested beam dump. Ramp-down and precycle.
  • 02h00: Fire alarm US15. Access.
  • 04h40: Access completed. Start recovery.
  • 05h27: Injecting pilot beams: large chroma errors. No beam from SPS for continuing. Pre-cycle.
  • 07h23: Injection: 3 x 1.1e11 p per beam.
  • 09h06: Start ramp.
  • 10h37: Start squeeze.
  • 11h56: Stable beams (fill 1185). New record luminosity of up to 8e29 cm-2 s-1.
  • 16h15: Beam dump during thunderstorm. Rampdown and precycle.
  • 17h45: Access in point 1 (QPS) and point 4 (wire scanner).
  • 19h00: End of access. Recovery and prepare for injection.
  • 20h44: Controls problems due to AFS failure.
  • Beam uptime: 10h30
  • Stable beam: 4h19
Report from 8h30 meeting

Monday June 28th

  • Recovery from cryo stop that happened Sunday, June 27th, 15h33.
  • 10h40: Access IR4 for wire scanner. Access in CMS. Access in 7-8 for BLM intervention.
  • 12h40: Tests of squeeze sequence without beam in simulation mode. Completed successfully.
  • 15h15: Cryo OK back.
  • 18h30: Preparation for injection completed (pre-cycle, RF, injection kickers, ...). Injection.
  • 19:00: Lost cryo in 3-4 (level regulation problem in DFBAF_O7R3). Waiting for cryo OK to start another precycle.
  • 19h46: Pre-cycle.
  • 22h30: Injection. Corrected bump in IR6 (MCBCH.8L6.B2 and MCBYH.5R6.B2).
  • 23h50: Fire alarm US15. Reset.
  • Beam uptime: 1h50
Report from 08:30 meeting